Darla is building a new desk. To make sure she had made a square corner, she measures 6 ft. from the corner along the edge and 4 ft. from the corner along the other edge. How long should the diagonal be between those two points if the corner is a right angle?
Answer by mananth(16946)

Darla is building a new desk. To make sure she had made a square corner, she measures 6 ft. from the corner along the edge and 4 ft. from the corner along the other edge. How long should the diagonal be between those two points if the corner is a right angle?
..
length = 6 feet
width = 4 feet
...
6^2+4^2=36+16=52
diagonal = sqrt52
diagonal = 7.21 feet